Use generating functions to solve the recurrence relation a_k=3a_(k-1)+4^(k-1) with the initial condition a_0=1.
1
Expert's answer
2020-06-08T21:13:53-0400

Given "a_0 = 1", and "a_k=3a_{k-1}+4^{k-1}" .

So, "a_1 = 3a_0 + 4^0 = 3 + 1 = 4"

"a_2 = 3a_1 + 4^1 = 12+4 = 16 \\
a_3 = 3a_2 + 4^2 = 48 + 16 = 64 \\
a_4 = 3a_3+4^3 = 192+64 = 256" and so on.

Hence the given series is "1, 4 , 16 , 64, 256" , _ _ _ _.

Hence, the generating function is "a_n = 4^{n}" .